# Quantitative Understanding of PDF Fits and their Uncertainties

### v1

@@ 5 Conclusions
Our present age is marked by unprecedented advancements in machine learning techniques, whose applications span many scientific domains – PDF determination being one of them. However, we believe that we ought to move from mere witnesses to active contributions. In fact, it is of paramount importance to understand how these new techniques behave when applied to complex problems such as PDF determination.
@@
Yet, this work is far from being conclusive. As a pioneering study, we believe that the most important contribution is the identification and initial discussion of the role of the NTK in PDF fits, leading to a set of relevant diagnostic metrics to be explored in future works. Indeed, many aspects merit further investigations. First of all, we need to study in detail the outcome of a similar analysis when applied to a more complex and realistic framework, including multiple fitted flavors and real data. For data that depend linearly on the PDFs, the extension of the formalism is straightforward. However, the agreement with the analytical predictions need to be tested by performing the numerical simulations. Further insights could be gained by exploring different parametrizations beyond neural networks. These systematic studies will be carried out in forthcoming works using the recent flexible and extensible open-source framework Colibri [Costantini:2025agd].
@@ Appendix A The BCDMS dataset for $T_{3}$
The analysis presented in this work employs a set of synthetic data points generated using a known underlying law $f^{\rm in}$ that we seek to reconstruct. Analogous to Ref. [Candido:2024hjt], the pseudo-data are constructed by convolving $f^{\rm in}$ with FK tables whose kinematic dependence is specified by measurements of the structure function $F_{2}$ on proton and deuterium targets from the BCDMS collaboration [Benvenuti:1989fm]. By combining these measurements, one can construct the observable $F_{2}^{p}-F_{2}^{d}$ that, at next-to-next-to-leading order (NNLO) in QCD and under the assumption of isoscalarity for the deuterium nucleus, provides a clean probe of the non-singlet triplet PDF combination $T_{3}=u^{+}-d^{+}$, where $u^{+}=u+\bar{u}$ and $d^{+}=d+\bar{d}$. The factorisation formula in Eq. (1) then simplifies to
@@
where $C_{T_{3}}$ is the corresponding Wilson coefficient computed in perturbative QCD, and $\otimes$ is a short-hand notation that denotes the convolution as in Eq. (1). The construction of the FK tables needed to compute the predictions, as well as the construction of the covariance matrix, are identical to Ref. [Candido:2024hjt], to which the reader is referred for further details. This yields a total of 333 data points, which then reduce to 248 points after applying the kinematic cuts.
Following the closure test framework developed by the NNPDF collaboration [DelDebbio:2021whr, NNPDF:2021njg], we generate data with three different levels of noise, labelled as Level 0 (L0), Level 1 (L1) and Level 2 (L2) data. Furthermore, we use the non-singlet triplet $xT_{3}$ from the NNPDF4.0 parton set [NNPDF:2021njg] as the input law $f^{\rm in}$. In the following, we summarise the definition of the different levels of pseudo-data.

### v2

@@ 5 Conclusions
Our present age is marked by unprecedented advancements in machine learning techniques, whose applications span many scientific domains – PDF determination being one of them. However, we believe that we ought to move from being mere witnesses to making active contributions. In fact, it is of paramount importance to understand how these new techniques behave when applied to complex problems such as PDF determination.
@@
Yet, this work is far from being conclusive. As a pioneering study, we believe that the most important contribution is the identification and initial discussion of the role of the NTK in PDF fits, leading to a set of relevant diagnostic metrics to be explored in future works. Indeed, many aspects merit further investigations. First of all, we need to study in detail the outcome of a similar analysis when applied to a more complex and realistic framework, including multiple fitted flavors and real data. For data that depend linearly on the PDFs, the extension of the formalism is straightforward. However, the agreement with the analytical predictions need to be tested by performing the numerical simulations. Further insights could be gained by exploring different parametrizations beyond neural networks. These systematic studies will be carried out in forthcoming works using the recent flexible and extensible open-source framework Colibri [18].
@@ Appendix A The BCDMS dataset for $T_{3}$
The analysis presented in this work employs a set of synthetic data points generated using a known underlying law $f^{\rm in}$ that we seek to reconstruct. Analogous to Ref. [14], the pseudo-data are constructed by convolving $f^{\rm in}$ with FK tables whose kinematic dependence is specified by measurements of the structure function $F_{2}$ on proton and deuterium targets from the BCDMS collaboration [13]. By combining these measurements, one can construct the observable $F_{2}^{p}-F_{2}^{d}$ that, at next-to-next-to-leading order (NNLO) in QCD and under the assumption of isoscalarity for the deuterium nucleus, provides a clean probe of the non-singlet triplet PDF combination $T_{3}=u^{+}-d^{+}$, where $u^{+}=u+\bar{u}$ and $d^{+}=d+\bar{d}$. The factorisation formula in Eq. (1) then simplifies to
@@
where $C_{T_{3}}$ is the corresponding Wilson coefficient computed in perturbative QCD, and $\otimes$ is a short-hand notation that denotes the convolution as in Eq. (1). The construction of the FK tables needed to compute the predictions, as well as the construction of the covariance matrix, are identical to Ref. [14], to which the reader is referred for further details. This yields a total of 333 data points, which then reduce to 248 points after applying the kinematic cuts.
Following the closure test framework developed by the NNPDF collaboration [21, 9], we generate data with three different levels of noise, labelled as Level 0 (L0), Level 1 (L1) and Level 2 (L2) data. Furthermore, we use the non-singlet triplet $xT_{3}$ from the NNPDF4.0 parton set [9] as the input law $f^{\rm in}$. In the following, we summarise the definition of the different levels of pseudo-data.